The measure of an angle is 21 degrees more than twice its supplement. Find the measure of each angle?
Answer by nerdybill(5404)

The measure of an angle is 21 degrees more than twice its supplement. Find the measure of each angle?
.
If two angles are supplementary, the sum of the two angles equals 180 degrees.
.
Let x = angle
then
180-x = supplement of angle
.
x = 2(180-x)+21
x = 360-2x+21
x = 381-2x
3x = 381
x = 127 degrees (angle)
.
supplement:
180-x = 180-127 = 53 degrees (supplement)